# Augmenting the action space with conventions to improve multi-agent cooperation in Hanabi

### v1

@@ Abstract
The card game Hanabi is considered a strong medium for the testing and development of multi-agent reinforcement learning (MARL) algorithms, due to its cooperative nature, hidden information, limited communication and remarkable complexity. Previous research efforts have explored the capabilities of MARL algorithms within Hanabi, focusing largely on advanced architecture design and algorithmic manipulations to achieve state-of-the-art performance for a various number of cooperators. However, this often leads to complex solution strategies with high computational cost and requiring large amounts of training data. For humans to solve the Hanabi game effectively, they require the use of conventions, which often allows for a means to implicitly convey ideas or knowledge based on a predefined, and mutually agreed upon, set of “rules”. Multi-agent problems containing partial observability, especially when limited communication is present, can benefit greatly from the use of implicit knowledge sharing. In this paper, we propose a novel approach to augmenting the action space using conventions, which act as special cooperative actions that span over multiple time steps and multiple agents, requiring agents to actively opt in for it to reach fruition. These conventions are based on existing human conventions, and result in a significant improvement on the performance of existing techniques for self-play and cross-play across a various number of cooperators within Hanabi.
@@ 1 Introduction
Reinforcement learning (RL) holds promising potential to address a large variety of problems where artificial agent operation offers a significant improvement over alternative methods, such as hard-coded algorithms or solutions [1]. For certain real-world problems, single-agent operation is not optimal (or even possible) and the incorporation of multiple agents would be beneficial (or necessary). Scenarios such as autonomous vehicles navigating terrain [2], guided drone swarms [3], mapping verbal instructions to executable actions [4], or the switching of railway lines benefit from the use of multiple agents. RL can help agents to effectively cooperate within these multi-agent scenarios by learning from past and or simulated experiences.
@@
Unfortunately, the introduction of multiple agents into an environment typically increases the complexity of the problem exponentially. It introduces a moving target learning problem [5], since all the agents must learn simultaneously. The reason for this is that each individual agent’s policy changes over time, which in turn causes a non-stationary environment. This often inhibits all the agents from developing effective policies and can lead to undesired behaviour [5].
@@
One of the most straightforward solutions to multi-agent reinforcement learning (MARL) is the combination of deep Q-networks (DQNs) and independent Q-learning [6], where each agent independently and simultaneously learns its own action-value function (or Q-values [^1] Action values ($Q$) refer to the value assigned to a certain action $a$ within a given state $s$.) while interacting with the same environment. However, this strategy does not fair well when paired with partial observability [^2] Partial observability refers to problems where the full state space is hidden from each individual agent, resulting in each agent having their own unique perspective of the problem. [7, 8]. Hausknecht and Stone [9] have shown that recurrent neural networks offer an improved solution to MARL problems containing partial observability. These networks incorporate a built-in short term memory over which experiences are unrolled (or combined) to form longer sequences. This is often combined with deep Q-learning’s feed forward neural network to produce deep recurrent Q-learning [9].
@@
Games are often used as a medium for testing and evaluating RL algorithms, since they incorporate real world problems with well-defined rules and a clear metric for measuring performance, where examples include Go [20], Backgammon [21], and Dota 2 [22]. Hanabi is a cooperative card game containing partial observability and limited communication, requiring players to logically reason over the intentions and actions of their cooperators, a concept known as theory of mind [23]. Human players require the development of conventions [^3] Philosopher David Lewis defines conventions as an arbitrary, self-perpetuating solution to a recurring coordination problem [24]. in order to solve the problem effectively. Conventions have evolved with humanity throughout the years and range from driving on a certain side of the road to social conventions or norms [24].
Most research efforts on the Hanabi problem use complex architecture design with advanced algorithms, often focusing on estimating and calculating the beliefs of other agents [25, 13]. However, one avenue of the Hanabi problem yet to be explored is that of conventions, and how to incorporate existing human conventions into MARL solution strategies. In this paper, we propose a method to incorporate human conventions into MARL through the use of artificial conventions, which act as cooperative actions that span over multiple time steps and agents, and show how it significantly improves on the performance of MARL agents within Hanabi for self-play [26] as well as cross-play [27] scenarios. Our approach shares similarities with options, due to the multi-time step extension of actions, however it is fundamentally built on a different concept.
@@ 1.1 Related Work
Hanabi was first proposed as a viable medium for MARL by Bard et al. [36], with focus placed on the philosophical ideas and challenges found within the problem setting and how they translate to real world scenarios. Bard et al. conducted initial tests for three different MARL algorithms, namely MA-Rainbow, actor-critic-Hanabi-agent (ACHA) [36], and Bayesian action decoder (BAD) [25], and compared them to state-of-the-art handcrafted bots, such as HatBot [37] and WTFWThat [38]. The MARL solutions achieved an acceptable score with the best performing agent achieving 16.8/25 for the most difficult scenario of five players, but still significantly falls behind the best handcrafted bot which achieves a score of 24.89/25.
@@
Hu et al. [27] also builds on the concept of the SAD algorithm, but rather than achieving state-of-the-art performance for self-play agents, they focus on the agent’s ability to cooperate with a variety of partners–a concept known as cross-play. This is achieved using other-play, an architecture for multi-agent systems specifically designed to break the ties formed between self-play agents, allowing them to produce more robust strategies. This allows the agents to cooperate with agents from different training runs (or regimes), and even human cooperators [27]. Even though the other-play agents are able to achieve effective cross-play performance when paired with a variety of agents in a two player scenario, the research does not focus on higher player counts where more advanced cooperation strategies are required. Additionally, other-play builds on the existing architecture of the SAD algorithm leading to increased computational complexity and cost [27].
@@ 1.2 Summary of Contributions
Rather than focusing on complex architecture design, this research shifts focus to incorporating existing domain knowledge of the Hanabi problem into MARL. Hanabi has an active and dedicated community with players constantly searching for new conventions and ways to reliably beat the game. These conventions are often based on principles, and have generally been standardised within the Hanabi community. Our main contribution is to show how these existing human conventions can be implemented in a MARL scenario using artificial conventions, which act as a special form of cooperative actions and action space augmentations.
@@
One of the key insights to this approach is our “subscribing” technique for convention continuation, where an agent can initiate a convention (based on a certain human convention) and subsequently a new action appears to the other agents to “subscribe” or continue with this convention. This strategy requires that agents opt in to a certain convention in order for it to reach completion, and allows an agent to halt a certain inferior convention, and/or initiate an improved convention, based on their unique observation. To this end, each convention contains an initial condition acting as the start of a convention, subsequent continuation conditions for other agents to opt in, a termination condition for a terminating action, and finally a policy that translates each aspect of the convention into an environment action based on the current agent’s observation.
@@
Our approach shares similarities with that of multi-agent options, but addresses a problem where agents must often wait for other agents to complete their options before they can initiate a new option (which can include a cooperative option) [46]. Fundamentally, our approach is built on a different concept, due to the cooperative requirement of conventions to reach fruition, and their ability to convey ideas or intentions through implicit communication based on mutually agreed upon “rules”. Therefore, it is important to note that our approach does not introduce additional explicit communication, or additional communication channels for that matter, but rather allows for implicit communication through the nature of conventions.
@@
Conventions result in a large performance uplift for existing MARL techniques, in our case multi-agent Rainbow, by significantly decreasing the training time and, in the case of 3–5 player Hanabi, increasing the converged performance. Additionally, since the agents are learning from a communal and well-defined list of conventions, the agents are not just able to cooperate in self-play, but also in a cross-play setting with agents from different training runs (or regimes). Cross-play has been shown to be an important research area for MARL since it opens the door to agents being able to cooperate with never-before-seen teammates, whether they are agents from other regimes, agents with different architectures, or humans [47, 48].
@@ 1.3 Paper Outline
In Section 2 we introduce the background and theory of RL, MARL, and options. In Section 3, we discuss the Hanabi environment as well as existing solution strategies, with focus placed on human conventions, in order to give context for our approach. This is followed by the discussion on conventions and action space augmentation with conventions in Section 4. Finally, we present and discuss our results for self-play and cross-play for all player counts in Hanabi, in Section 5.
@@ 2 Multi-agent Reinforcement Learning and Options
where $S$ denotes the global state space and $A$ represents the joint-action space of the $n$ agents at time $t$. The observation space $O$ consists of the local observation of each agent $i$ at time $t$ within the global state $S$ ($O^{i}_{t}=O(S;i;t)$). After $A$ is sent to the environment, the global state transitions from $S$ to $S^{\prime}$ given the state transition probability $T$, and similarly $P$ represents the observation transition probability to transition from $O$ to $O^{\prime}$. This results in the environment producing $R$, which donates the immediate reward for each agent action that contributed to a global state change at $t$, with $\gamma$ denoting the discount factor for future reward. These rewards are either combined with equal weighting (summed), or can have their own discount factor $\gamma_{r}$, and is referred to as the forward accumulated reward [50].
@@ 2.1 Independent Q-learning
where $\alpha$ is the learning rate and $\gamma$ the discount factor of future rewards [49]. This can be extended to MARL with independent Q-learning, where each agent has their own action-value function and uses their individual experiences in their update steps [51]. The pseudocode for independent Q-learning in a turn-based scenario is shown in Algorithm 1, where each agent $i$ uses their own observations $O^{i}_{t}$ to update their action-value functions $Q^{i}$. Note that due to the unique nature of turn-based settings, a special type of forward accumulated reward (FAR) must be used which consists of the one round return resulting from each agent’s action, i.e. $\sum_{k=0}^{P-1}R_{t+1+k}$.
@@
10:              $\sum_{k=0}^{P-1}R_{t+1+k}$triangleright Calculate the FAR based on the rewards for the next round of actions resulting from $A^{i}_{t}$
11:              $Q^{i}(O^{i}_{t},A^{i}_{t})\leftarrow Q^{i}(O^{i}_{t},A^{i}_{t})+\alpha\bigr{[}\sum_{k=0}^{P-1}R_{t+1+k}+\gamma\max_{a}Q^{i}(O^{i}_{t+P},a)-Q^{i}(O^{i}_{t},A^{i}_{t})\bigr{]}$
@@
Instead of distinct policies, independent Q-learning agents can also make use of a shared policy, especially when the environment contains symmetries [51]. This allows the update step (2) to be updated more frequently by using the experiences of all the independent agents. This has proven to offer significant performance gain and allow agents to learn more effectively [51].

### v2

@@ Abstract
The card game Hanabi is considered a strong medium for the testing and development of multi-agent reinforcement learning (MARL) algorithms, due to its cooperative nature, hidden information, limited communication and remarkable complexity. Previous research efforts have explored the capabilities of MARL algorithms within Hanabi, focusing largely on advanced architecture design and algorithmic manipulations to achieve state-of-the-art performance for various number of cooperators. However, this often leads to complex solution strategies with high computational cost and requiring large amounts of training data. For humans to solve the Hanabi game effectively, they require the use of conventions, which often allows for a means to implicitly convey ideas or knowledge based on a predefined, and mutually agreed upon, set of “rules” or principles. Multi-agent problems containing partial observability, especially when limited communication is present, can benefit greatly from the use of implicit knowledge sharing. In this paper, we propose a novel approach to augmenting an agent’s action space using conventions, which act as a sequence of special cooperative actions that span over and include multiple time steps and multiple agents, requiring agents to actively opt in for it to reach fruition. These conventions are based on existing human conventions, and result in a significant improvement on the performance of existing techniques for self-play and cross-play across various number of cooperators within Hanabi.
@@ 1 Introduction
Reinforcement learning (RL) holds promising potential to address a large variety of problems where artificial agent operation offers a significant improvement over alternative methods, such as hard-coded algorithms or solutions [1]. For certain real-world problems, single-agent operation is not optimal (or even possible) and the incorporation of multiple agents would be beneficial (or necessary). A prime example is autonomous vehicles navigating roads [2], where each vehicle is controlled by an autonomous agent and these agents must cooperate effectively to ensure road safety. Other examples of problems that benefit from multiple agents include guided drone swarms [3], mapping verbal instructions to executable actions [4], or the switching of railway lines. RL can help agents to effectively cooperate within these multi-agent scenarios by learning from past and or simulated experiences.
@@
Unfortunately, the introduction of multiple agents into an environment typically increases the complexity of the problem exponentially. It introduces a moving target learning problem [5], since all the agents must learn simultaneously. The reason for this is that each individual agent’s policy changes over time, which in turn causes a non-stationary environment. This often inhibits all the agents from developing effective policies and can lead to undesired behaviour [5]. Furthermore, multi-agent systems often contain partial observability, where the full state space is hidden from the individual agents, resulting in each agent having their own unique perspective of the problem. In the case of autonomous agents controlling vehicles, even though an agent might have access to high-definition sensors (often used in vehicle control), the environment will still contain objects outside the current agent’s perspective, for example objects outside the sensor’s range or viewing angle or objects obscured by other vehicles, static objects, blind corners, concealed driveways, etc.
When another agent controlling a different vehicle is introduced into the environment, the need for communication to overcome the problem of partial observability becomes apparent, since this agent will have a different viewpoint of the environment. These viewpoints will have some overlap, but will also contain important information that the agents will have to share, such as obscured objects, accidents, obstacles, pedestrians, etc., in order to safely navigate the road. Even when the agents are able to communicate, it is infeasible to share and process all the high quality information between agents, especially in scenarios where quick reaction times are crucial. This communication problem becomes even more difficult if we consider situations where more than two vehicles must coordinate effectively, for example a large intersection, road construction, an accident that lead to congestion, or a busy city.
@@
One of the most straightforward solutions to multi-agent reinforcement learning (MARL) is the combination of deep Q-networks (DQNs) and independent Q-learning [6], where each agent independently and simultaneously learns its own action-value function (or Q-values [^1] Action values ($Q$) refer to the value assigned to a certain action $a$ within a given state $s$.) while interacting with the same environment. However, this strategy does not fair well when paired with partial observability [7, 8]. Hausknecht and Stone [9] have shown that recurrent neural networks offer an improved solution to MARL problems containing partial observability. These networks incorporate a built-in short term memory over which experiences are unrolled (or combined) to form longer sequences. This is often combined with deep Q-learning’s feed forward neural network to produce deep recurrent Q-learning [9].
@@
Games are often used as a medium for testing and evaluating RL algorithms, since they incorporate real-world problems with well-defined rules and a clear metric for measuring performance, where examples include Go [20], Backgammon [21], and Dota 2 [22]. Hanabi is a cooperative card game containing partial observability and limited communication, requiring players to logically reason over the intentions and actions of their cooperators, a concept known as theory of mind [23]. Human players require the development of conventions [^2] Philosopher David Lewis defines conventions as an arbitrary, self-perpetuating solution to a recurring coordination problem [24]. in order to solve the problem effectively. Conventions have evolved with humanity throughout the years and range from driving on a certain side of the road to social conventions or norms [24].
Conventions allow for additional implicit communication through actions and mutually agreed upon “rules”, thereby overcoming the communication restrictions of the communication channel. Similar to the Hanabi problem, when humans tackle the problem of controlling vehicles on roads, conventions are common practice and ensure effective cooperation. Examples include: driving on a certain side of the road [^3] Philosopher David Lewis discusses this topic in depth, and places particular focus on how this affects other aspects of life, such as people tending to keep to the same side as their road conventions when walking in a busy street [24].; keep left, pass right (or vice versa); who has rite of passage at a T-intersection or four-way intersection (and how this changes based on the presence of pedestrians); flashing of headlights while stationary to indicate giving right of way/yielding; flashing of headlights while moving to indicate danger ahead, or be wary of potential danger ahead; flashing of hazards to indicate an emergency; a quick flash of hazards to say thank you; and many more. Furthermore, this highlights the importance for artificial agents to learn existing human conventions, since there will exist a period of time when autonomous agents must navigate these complex scenarios alongside humans, and the agents will have to be able to convey and receive information effectively without using their agent-specific communication channel.
Most research efforts on the Hanabi problem use complex architecture design with advanced algorithms, often focusing on estimating and calculating the beliefs of other agents [25, 13]. However, one avenue of the Hanabi problem yet to be explored is that of conventions, and how to incorporate existing human conventions into MARL solution strategies. In this paper, we propose a method to incorporate human conventions into MARL through the use of artificial conventions, which acts as cooperative actions that span over and include multiple time steps and multiple agents, and show how it significantly improves on the performance of MARL agents within Hanabi for self-play [26] as well as cross-play [27] scenarios. Our approach shares similarities with options, due to the multi-time step extension of actions, however is fundamentally built on a different concept.
@@ 1.1 Related Work
Hanabi was first proposed as a viable medium and frontier for MARL by Bard et al. [36], with focus placed on the philosophical ideas and challenges found within the problem setting and how they translate to real-world scenarios. Bard et al. conducted initial tests for three different MARL algorithms, namely MA-Rainbow, actor-critic-Hanabi-agent (ACHA) [36], and Bayesian action decoder (BAD) [25], and compared them to state-of-the-art handcrafted bots, such as HatBot [37] and WTFWThat [38]. In the most difficult scenario of five players, the MARL solutions achieved an acceptable score of 16.8/25, however this still significantly falls behind the best handcrafted bot able to achieve a score of 24.89/25.
@@
Hu et al. [27] also builds on the concept of the SAD algorithm, but rather than achieving state-of-the-art performance for self-play agents, they focus on the agent’s ability to cooperate with a variety of partners–a concept known as cross-play. This is achieved using other-play, an architecture for multi-agent systems specifically designed to break the ties formed between self-play agents, allowing them to produce more robust strategies. This allows the agents to cooperate with agents from different training runs (or regimes), and even human cooperators [27]. Even though the other-play agents are able to achieve effective cross-play performance when paired with a variety of agents in a two player scenario, the research does not focus on higher player counts where more advanced cooperative strategies are required. Additionally, other-play builds on the existing architecture of the SAD algorithm leading to increased computational complexity and cost [27].
@@ 1.2 Summary of Contributions
Rather than focusing on complex architecture design, this research shifts focus to incorporating existing domain knowledge of the Hanabi problem into MARL algorithms. Hanabi has an active and dedicated community with players constantly searching for new conventions and ways to reliably beat the game. These conventions are often based on principles [^4] A convention principle is a strategy, or user defined “rule”, external to the existing game rules which governs a player’s behaviour and results in a desired outcome., and have generally been standardised within the Hanabi community. Our main contribution is to show how these existing human conventions can be implemented in a MARL scenario using artificial conventions, which act as a special form of cooperative actions and can be incorporated into existing MARL algorithms using action space augmentation.
@@
One of the key insights to our approach is the “subscribing” technique for convention continuation, where an agent can initiate a convention (based on a certain human convention) and subsequently a new action appears to the other agents to “subscribe” or continue with this convention. This strategy requires that agents opt in to a certain convention in order for it to reach completion, and allows an agent to halt a certain inferior convention, and/or initiate an improved convention, based on their unique observation. To this end, each convention contains an initial condition acting as the start of a convention, subsequent continuation conditions for other agents to opt in, a termination condition for a terminating action, and finally a policy that translates each step of the convention into an environment action based on the current agent’s observation.
@@
Our approach shares similarities with that of multi-agent options, but addresses a problem where agents must often wait for other agents to complete their options before they can initiate a new option (which can include a cooperative option) [46]. Fundamentally, our approach is built on a different concept, due to the cooperative requirement of conventions to reach fruition, and their ability to convey ideas or intentions through implicit communication based on mutually agreed upon “rules”. Therefore, it is important to note that our approach does not introduce additional explicit communication, or additional communication channels for that matter, but rather allows for implicit communication through the nature of conventions. Furthermore, each agent forming part of the convention actively participates in every step of the convention, rather than an external policy taking control and executing over multiple time steps until the option terminates.
Ultimately, we believe our work will act as the foundation for convention discovery, i.e. the ability for agents to define their conventions as they train. Similar to how options paved the way for option discovery by showing the importance and benefits of options, conventions offer a similar argument for multi-agent cooperation in partially observable environments. Option discovery has been proven to be a difficult problem to solve in multi-agent systems [46, 47], and we believe convention discovery would offer a similar challenge. Therefore, it is important to first prove the capabilities of conventions, their implications, implementation, and limitations, especially at the hands of a difficult problem such as Hanabi.
@@
Conventions result in a large performance uplift for existing MARL techniques, specifically multi-agent Rainbow, by significantly decreasing the training time and, in the case of 3–5 player Hanabi, increasing the converged performance. Additionally, since the agents are learning from a communal and well-defined list of principles, the agents are not just able to cooperate in self-play, but also in a cross-play setting with agents from different training runs (or regimes). Cross-play has been shown to be an important research area for MARL since it opens the door to agents being able to cooperate with never-before-seen teammates, whether they are agents from other regimes, agents with different architectures, or humans (such as autonomous agents navigating the road alongside humans) [48, 49].
@@ 1.3 Paper Outline
In Section 2 we introduce the background and theory of RL, MARL, and options, in order to give context for our approach. This is followed by the introduction and discussion of conventions and action space augmentation with conventions in Section 3. In Section 4, we introduce the Hanabi problem and learning environment, as well as existing human conventions and how they translate to conventions using our discussed formulations and definitions. Section 4 also contains the results for initial tests conducted on the Small Hanabi environment, to prove the capabilities and benefits of using action space augmentation with conventions. Finally, we present and discuss our results for self-play and cross-play for all player counts in Hanabi, in Section 5.
@@ 2 Multi-agent Reinforcement Learning and Options
where $S$ denotes the global state space and $A$ represents the joint-action space of the $n$ agents at time $t$. The observation space $O$ consists of the local observation of each agent $i$ at time $t$ within the global state $S$ ($O^{i}_{t}=O(S;i;t)$). After $A$ is sent to the environment, the global state transitions from $S$ to $S^{\prime}$ given the state transition probability $T$, and similarly $P$ represents the observation transition probability to transition from $O$ to $O^{\prime}$. This results in the environment producing $R$, which denotes the immediate reward for each agent action that contributed to a global state change at $t$, with $\gamma$ denoting the discount factor for future reward. These rewards are either combined with equal weighting (summed), or can have their own discount factor $\gamma_{r}$, and is referred to as the forward accumulated reward [51].
@@ 2.1 Independent Q-learning
where $\alpha$ is the learning rate and $\gamma$ the discount factor of future rewards [50]. This can be extended to MARL with independent Q-learning, where each agent has their own action-value function, conditioned on their observations $O^{i}_{t}$ instead of the global state $S_{t}$, and using their individual experiences in their update steps [52]. The algorithm for independent Q-learning in a turn-based scenario is shown in Algorithm 1, where each agent $i$ uses their own observations $O^{i}_{t}$ to update their action-value functions $Q^{i}$. The algorithm receives as input all the hyperparameters for the agent’s architecture, and produces a policy as output in the form of the action-value function.
@@
10:              $\sum_{f=0}^{P-1}R_{t+1+f}$triangleright Calculate the FAR based on the rewards for the next round of actions resulting from $A^{i}_{t}$
11:              $Q^{i}(O^{i}_{t},A^{i}_{t})\leftarrow Q^{i}(O^{i}_{t},A^{i}_{t})+\alpha\bigr{[}\sum_{f=0}^{P-1}R_{t+1+f}+\gamma\max_{a}Q^{i}(O^{i}_{t+P},a)-Q^{i}(O^{i}_{t},A^{i}_{t})\bigr{]}$
@@
Due to the turn-based nature of the environment, for each episode, the algorithm steps through each agent (line 6) until the global terminal state $S_{t+1}$ is reached. Since the algorithm depicts Q-learning, the agents use an $argmax$ function to select an action based on the Q-values for a given observation, shown in line 7. After an action has been chosen, it is sent to the environment, which reacts accordingly and produces a new observation and reward. Note that due to the unique nature of turn-based settings, a special type of forward accumulated reward (FAR) must be used which consists of the one round return resulting from each agent’s action [40], i.e. $\sum_{f=0}^{P-1}R_{t+1+f}$. Finally, the original observation, next observation, as well as FAR are used in line 11 to update the action-value function according to equation (2).
Instead of distinct policies $Q^{i}$, independent Q-learning agents can make use of a shared policy $Q$, especially when the environment contains symmetries [52]. This allows the update step in equation (2) to be updated more frequently by using the experiences of all the independent agents. This has proven to offer significant performance gain and allow agents to learn more effectively [52].